I'm in 7th grade. I am having problems dividing integers.
2 1/4x 2 1/9 =
I got
19/4 - 19/9 = 19/4 = 4 3/4

Answer by Alan3354(69443)

I'm in 7th grade. I am having problems dividing integers.
2 1/4x 2 1/9 =
This is not a division problem, and it's not integers.
I got
19/4 - 19/9 = 19/4 = 4 3/4 *** I don't know what you did, or how you got 19/4.
======================
======================
2 1/4x 2 1/9 =
-------
Change to improper fractions.
= (9/4)*(19/9) (use * for multiply, not x)
Multiply the NUMs and the DENs
= 9*19/(4*9)
You can cancel the 9 in the NUM and the DEN
= 19/4